# AI-Native 6G Physical Layer with Cross-Module Optimization and Cooperative Control Agents

### v1

@@ III-A Problem Formulation
As illustrated in the Fig. 1, the proposed AI-native framework includes the data flow and control strategy. For downlink data flow, it requires multiple candidate transmission schemes consisting of AI/ML features, e.g., AI/ML based modulation, AI/ML based precoding, AI/ML based CSI feedback and so on. Specifically, different compositions of the schemes constitutes different physical layer links $\mathfrak{F}$, i.e.,
@@
where $\mathfrak{M}_{q}(\cdot)$, $1\leq q\leq Q$, $\mathfrak{P}_{v}(\cdot)$, $1\leq v\leq V$ and $\mathfrak{C}_{z}(\cdot)$, $1\leq z\leq Z$ denote the utilized modulation, precoding and CSI feedback schemes with number of candidates model architectures $Q$, $V$ and $Z$, respectively. Specifically, $\Theta_{\mathrm{mod}}$ and $\Theta_{\mathrm{demod}}$ represent the model parameters of the modulation and demodulation part of $\mathfrak{M}_{q}(\cdot)$, respectively. $\Theta_{\mathrm{P}}$ denotes the model parameters of $\mathfrak{P}_{v}(\cdot)$. $\Theta_{\mathrm{E}}$ and $\Theta_{\mathrm{D}}$ denote the model parameters of the encoder and decoder part of $\mathfrak{C}_{z}(\cdot)$, respectively. $\mathfrak{O}$ denotes the other features that are set as default in this paper. Moreover, in order to leverage the potential of these AI/ML features, targeted enhancement mechanisms for each one and cross-module optimization are also proposed, which will be described in detail later.
@@ III-B1 Constellation Solution Space Expansion for Downlink Transmission
where $g_{\mathrm{demod}}(\cdot)$ represents the proposed AI/ML-based cross-layer demodulator with trainable model weights of $\Theta_{\mathrm{demod}}$, $\tilde{\mathbf{s}}^{\mathrm{eq}}_{f,t}\in\mathbb{C}^{N_{\mathrm{layer}}\times 1}$ and $\hat{\mathbf{c}}_{f,t,n_{\mathrm{layer}}}\in\{0,1\}^{m\times 1}$ with $n_{\mathrm{layer}}=1,\dots,N_{\mathrm{layer}}$ denote the equalized signal and received LLRs of subcarrier $f$ and OFDM symbol $t$, respectively. Within the proposed method, $g_{\mathrm{mod}}(\cdot)$ and $g_{\mathrm{demod}}(\cdot)$ directly learn optimal bits-to-symbols mapping, breaking the dimensional isolation between layers in conventional methods. This approach enables joint exploitation of cross-layer features and expands the solution space for optimal constellation search, thereby facilitating the learning of an better constellation set to further enhance modulation performance.
@@ III-B2 Utility-Oriented Precoder Construction
To address the above issues, a utility-oriented precoder construction method is proposed, where the decoder function is redefined by shifting its role from recovering the original CSI to directly generating precoding matrices optimized for end-to-end performance, using binary cross entropy error (BCE) loss function of
@@ III-B3 Bit-Level Bottlenecks Bypass for Uplink CSI Feedback
where $\hat{\Theta}_{\rm{E}}$ and $\hat{\Theta}_{\rm{D}}$ denote the trainable model weights of encoder and decoder respectively, $\mathbf{s}_{\mathrm{CSI}}\in\mathbb{C}^{N_{\mathrm{CSI}}\times 1}$ denotes the symbols for transmission in MIMO-OFDM system, $\hat{\mathbf{s}}\in\mathbb{C}^{N_{\mathrm{CSI}}\times 1}$ denotes the received symbols after channel equalization, and $N_{\mathrm{CSI}}=N_{\mathrm{CSIcoded}}/m$ with $m$ denoting the modulation order. Notably, the transition of the autoencoder’s intermediate variable from bit-form $\mathbf{c}_{\mathrm{CSI}}$ to non-bit-form $\mathbf{s}_{\mathrm{CSI}}$ eliminates quantization errors and training difficulties, where the continuous solution space permits gradient-based training to converge to local optima without being trapped by quantization boundaries, and ii) The preserved differentiability throughout the system allows joint training of constellation geometry with other modules, aligning the design with global performance objectives rather than isolated CSI reconstruction metrics. While absolute optimality cannot be guaranteed due to the non-convex nature of the problem, this expanded solution space provides necessary conditions for finding superior solutions, as empirically verified in the simulation results of subsequent Section IV.
@@ III-B4 Control Agent for Optimized Transmission Schemes
In this subsection, the design for control agent $\mathfrak{G}$ is introduced. From the perspective of system gain, the traditional link adaptation mechanism utilizes a lookup table mapping CQI to MCS, which is based on the BLER threshold analyzed from traditional non-AI-native links. However, once AI/ML features are introduced, the mapping rules are no longer easy to obtain by theoretical derivation due to the uninterpretability of AI/ML models. At the same time, the implementation form of link adaptation also changes from MCS selection to model switching. Existing solutions of LCM [43] adopt post-hoc adaptation, as the switch is conducted after performance degradation is detected during model monitoring. These call for novel design of control strategy for the optimized transmission scheme.
Here we propose an AI/ML-based strategy that directly learns transmission schemes selection using the data collected from the proposed AI-native link and target channel scenario. First, there is a one-to-one correspondence between the utility-oriented precoder and the CSI input to the encoder, where the utility-oriented precoder and the cross-layer-modulated symbols can adjust the transmission power of each layer, and indirectly adjust the number of transmission layers, e.g., the layers with zero power are not sent. Therefore, the proposed CSI feedback not only directly adjusts the precoder, but also indirectly adjusts the number of transmission layers. Second, AI/ML-based control agents are also proposed that learn the mapping between estimated channel quality of signal to interference and noise ratio (SINR), and the transmission schemes that optimize the throughput of AI-native link, i.e.,
@@ III-B5 Cross-Module Optimization Process
Phase II then deactivates the above auxiliary loss with $\lambda=1$, focusing exclusively on BCE-driven end-to-end refinement to finally achieve the proposed cross-layer modulation and utility-oriented precoder. Notably, channel encoding and decoding modules are excluded during training phase to ease the difficulty of training while being cascaded only in the inference phase to maintain compatibility with standard communication pipelines. It should be noted that the proposed bit-level bottlenecks bypass solution maintains the complete link differentiable for the above gradient-based optimization.
As for Phase III, the switching model $g_{\mathrm{la}}(\cdot)$ is trained using the loss of categorical cross entropy, i.e.,
@@ III-B6 AI/ML Model Structure Implementation
The model structure supporting the proposed CSI feedback in III-B2 and III-B3 is detailed in Table III. The Transformer backbone for CSI feedback namely EVCsiNet-T [5] with embedding dimension of $N_{\mathrm{embedding}}=256$, $N_{\mathrm{head}}=4$ heads and $N_{\mathrm{transformer}}=6$ basic blocks is implemented for feature extraction, wherein the EVCsiNet-T is a common model structure used by 3GPP to evaluate the performance of CSI feedback. Specifically, the quantization and dequantization layers are replaced by Dense layers with normalization that directly output normalized modulation constellation symbols.
@@ III-C Theoretical analysis
In this subsection, we take the module of multi-layer modulation as an instance, shedding light on the potential benefits for performing cross-layer modulation with theoretical analysis. In our proposed transmission scheme, the bit-interleaved coded modulation (BICM) design principle is still inherited, where the performance of modulation can be measured by BICM capacity [44]
@@
where $b\in\{0,1\}$ is equiprobable bit, $m$ denotes modulation order, ${\bf{y}},{\bf{x}}\in{\text{C}}^{N_{\rm{layer}}}$ denote the received constellation and transmitted constellation, respectively. Notation $\mathcal{X}_{b}^{i}$ represents the set of symbols where the $i$-th bit equals $b$ and $\mathcal{X}$ denotes the set of all symbols. For traditional modulation method, both $\bf{y}$} and $\bf{x}$} are complex scalar. On the contrary, $\bf{y}$} and $\bf{x}$} can be complex vectors. The expectation operator shall jointly consider the high-dimensional distribution of transmitted symbols and noise, which is hard to be directly analyzed.
@@
Due to the intractability of the optimization on $l_{b}^{(i)}({\bf{y_{s}}})$ for all symbols, we turn to the analysis of upper bound of error probability for each symbol, which indirectly affects channel capacity. Reviewing the exponential component in (25), the minimization of probability is equivalent to maximizing the Euclidean distance between adjacent constellations, i.e. maximizing the minimum distance $d_{\rm{min}}$. Intuitively, a lager $d_{\rm{min}}$ for the given dimension can effectively reduce the error probability and contributes to improved channel capacity.
@@
where $r$ denotes the amplitude of the considered point, $V(\text{C}^{N_{\rm{layer}}})$ denotes the volume of the whole space. It can be expected that the shaped of sphere could maximize the space utilization efficiency, resulting in the averaged power of
@@
where the last approximation is obtained with $N!\simeq(N/e)^{N}$. This results are consistent with the ultimate shaping gain in [46]. To summarize, it can be seen that enlarging the modulation dimension can lead to a improved error probability and implicitly higher BICM capacity compared with legacy QAM constellation.
@@ IV Simulation Results
Dwonlink Carrier frequency 3.5GHz
@@
To demonstrate the effectiveness of the proposed cross-module optimization (CMO) scheme, we evaluate two training variants, i.e., CMO I and CMO II, against several 5G-compliant baselines under identical system constraints. CMO I exclusively undergoes phase I training until convergence, while CMO II sequentially executes phase I followed by phase II training. Both CMO I and CMO II share identical model implementations and total training iterations but employ distinct training processes. The conventional comparative scheme, denoted as 5G baselines, utilize standardized constellations for modulation and demodulation, including quadrature phase shift keying (QPSK), 16 quadrature amplitude modulation (QAM), 64QAM, and 256QAM, and supports a configurable number of transmission layers ranging in $N_{\mathrm{layer}}\in\{1,2,3,4\}$. The eTypeII codebook-based CSI feedback is adopted where eigenvectors that are fed back at the BS serve as the precoding matrix. To comprehensively assess baselines performance, three classical detection algorithms are employed. The LMMSE receiver aims to minimize the mean square error, thereby balancing interference suppression with noise enhancement. In contrast, the Zero-Forcing (ZF) receiver focuses solely on complete interference cancellation, which often leads to significant noise amplification. As a non-linear alternative, the K-Best receiver, a sub-optimal sphere decoding algorithm, provides a favorable performance-complexity trade-off by approximating the Maximum Likelihood (ML) performance. To focus solely on the downlink data transmission performance gains of the proposed CMO schemes, idealized uplink transmission conditions are assumed first. The CSI encoder and decoder components of both CMO I and CMO II retain the original EVCsiNet-T [5] architecture to ensure compatibility with CSI feedback overhead in bitstream form.
@@
As depicted in Fig. 4a to Fig. 4e, the BLER performance across varying downlink payload sizes is compared between the CMO II and CMO III approaches. The CMO III demonstrates consistent gains across all payload configurations, where the gain increasing as the payload size grows. Notably, the CMO III scheme achieves more pronounced advantages under poorer uplink signal conditions. At an uplink SNR of -10 dB, taking BLER = 0.1 as a reference, the performance gain over the CMO II grows from approximately 1 dB in low-bit regions to nearly 5 dB in high-bit regions. In more severe scenarios with uplink SNR of -20 dB, the disparity widens significantly. That is, the gain escalates from around 4 dB at 2 bits per RE to up to 10 dB at 32 bits per RE under the same BLER threshold of 0.1. These results clearly demonstrate that the design enhancement for CSI feedback in CMO III effectively boosts communication performance in scenarios with limited uplink link quality. As illustrated in Fig. 4f, the throughput performance of both CMO II and CMO III under CCA is compared across varying signal conditions. Consistent with the trends observed in the BLER results, the CMO III exhibits more substantial throughput gains at lower uplink SNR, highlighting its robustness in UE power-constrained situation. At an uplink SNR of -10 dB, the CMO III achieves a throughput gain of approximately 22% over the CMO II when the downlink SNR is 10 dB. In contrast, under a more challenging uplink SNR of -20 dB, the gain increases to around 52% at the same downlink SNR level. These further explain the advantages of the proposed solutions.
@@
To rigorously validate the robustness of the proposed CMO III under more realistic and challenging conditions, we extend the evaluation to a scenario involving two BSs and two UEs coexisting. For downlink data transmission, the two BSs transmit signals over the same time-frequency resources, leading to mutual inter-BS interference. For uplink feedback, the two UEs also share identical resources for signal transmission, resulting in mutual inter-UE interference. In this experimental setup, the target-to-interference power ratio is set to 1/0.3 to mimic practical interference conditions. As illustrated in Fig. 6, under a payload size of 8 bits per RE, CMO III still maintains a significant performance gain over CMO II. Notably, the gain of CMO III over CMO II is more pronounced at a UL SNR of -20 dB than at -10 dB, aligning with the general trend observed in the interference-free scenario. Comparing Fig. 6 with the interference-free scenario in Fig. 4 reveals that CMO III exhibits minimal performance degradation, significantly smaller than that of CMO II. This result confirms that the end-to-end training under interference conditions enhances the overall interference tolerance at both the transmitter and receiver ends. And more specifically, the enhanced CSI feedback mechanism in CMO III provides superior robustness against uplink interference, making it less affected by such interference than CMO II.

### v2

@@ III-A Problem Formulation
As illustrated in the Fig. 1, the proposed AI-native framework includes the data flow and control strategy. For downlink data flow, it requires multiple candidate transmission schemes consisting of AI/ML features, e.g., AI/ML based modulation, AI/ML based precoding, AI/ML based CSI feedback and so on. Specifically, different compositions of the schemes constitute different physical layer links $\mathfrak{F}$, i.e.,
@@
where $\mathfrak{M}_{q}(\cdot)$, $1\leq q\leq Q$, $\mathfrak{P}_{v}(\cdot)$, $1\leq v\leq V$ and $\mathfrak{C}_{z}(\cdot)$, $1\leq z\leq Z$ denote the utilized modulation, precoding and CSI feedback schemes with number of candidate model architectures $Q$, $V$ and $Z$, respectively. Specifically, $\Theta_{\mathrm{mod}}$ and $\Theta_{\mathrm{demod}}$ represent the model parameters of the modulation and demodulation part of $\mathfrak{M}_{q}(\cdot)$, respectively. $\Theta_{\mathrm{P}}$ denotes the model parameters of $\mathfrak{P}_{v}(\cdot)$. $\Theta_{\mathrm{E}}$ and $\Theta_{\mathrm{D}}$ denote the model parameters of the encoder and decoder part of $\mathfrak{C}_{z}(\cdot)$, respectively. $\mathfrak{O}$ denotes the other features that are set as default in this paper. Moreover, in order to leverage the potential of these AI/ML features, targeted enhancement mechanisms for each one and cross-module optimization are also proposed, which will be described in detail later.
@@ III-B1 Constellation Solution Space Expansion for Downlink Transmission
where $g_{\mathrm{demod}}(\cdot)$ represents the proposed AI/ML-based cross-layer demodulator with trainable model weights of $\Theta_{\mathrm{demod}}$, $\tilde{\mathbf{s}}^{\mathrm{eq}}_{f,t}\in\mathbb{C}^{N_{\mathrm{layer}}\times 1}$ and $\hat{\mathbf{c}}_{f,t,n_{\mathrm{layer}}}\in\{0,1\}^{m\times 1}$ with $n_{\mathrm{layer}}=1,\dots,N_{\mathrm{layer}}$ denote the equalized signal and received LLRs of subcarrier $f$ and OFDM symbol $t$, respectively. Within the proposed method, $g_{\mathrm{mod}}(\cdot)$ and $g_{\mathrm{demod}}(\cdot)$ directly learn optimal bits-to-symbols mapping, breaking the dimensional isolation between layers in conventional methods. This approach enables joint exploitation of cross-layer features and expands the solution space for optimal constellation search, thereby facilitating the learning of a better constellation set to further enhance modulation performance.
@@ III-B2 Utility-Oriented Precoder Construction
To address the above issues, a utility-oriented precoder construction method is proposed, where the decoder function is redefined by shifting its role from recovering the original CSI to directly generating precoding matrices optimized for end-to-end performance, using binary cross-entropy error (BCE) loss function of
@@ III-B3 Bit-Level Bottlenecks Bypass for Uplink CSI Feedback
where $\hat{\Theta}_{\rm{E}}$ and $\hat{\Theta}_{\rm{D}}$ denote the trainable model weights of encoder and decoder respectively, $\mathbf{s}_{\mathrm{CSI}}\in\mathbb{C}^{N_{\mathrm{CSI}}\times 1}$ denotes the symbols for transmission in MIMO-OFDM system, $\hat{\mathbf{s}}\in\mathbb{C}^{N_{\mathrm{CSI}}\times 1}$ denotes the received symbols after channel equalization, and $N_{\mathrm{CSI}}=N_{\mathrm{CSIcoded}}/m$ with $m$ denoting the modulation order. Notably, the transition of the autoencoder’s intermediate variable from bit-form $\mathbf{c}_{\mathrm{CSI}}$ to non-bit-form $\mathbf{s}_{\mathrm{CSI}}$ eliminates quantization errors and training difficulties, where the continuous solution space permits gradient-based training to converge to local optima without being trapped by quantization boundaries, and the preserved differentiability throughout the system allows joint training of constellation geometry with other modules, aligning the design with global performance objectives rather than isolated CSI reconstruction metrics. While absolute optimality cannot be guaranteed due to the non-convex nature of the problem, this expanded solution space provides necessary conditions for finding superior solutions, as empirically verified in the simulation results of subsequent Section IV.
@@ III-B4 Control Agent for Optimized Transmission Schemes
In this subsection, the design for control agent $\mathfrak{G}$ is introduced. From the perspective of system gain, the traditional link adaptation mechanism utilizes a lookup table mapping CQI to MCS, which is based on the BLER threshold analyzed from traditional non-AI-native links. However, once AI/ML features are introduced, the mapping rules are no longer easy to obtain by theoretical derivation due to the uninterpretability of AI/ML models. At the same time, the implementation form of link adaptation also changes from MCS selection to model switching. Existing solutions of LCM chen20235g adopt post-hoc adaptation, as the switch is conducted after performance degradation is detected during model monitoring. These call for novel design of control strategy for the optimized transmission scheme.
Here we propose an AI/ML-based strategy that directly learns the transmission schemes selection using the data collected from the proposed AI-native link and target channel scenario. First, there is a one-to-one correspondence between the utility-oriented precoder and the CSI input to the encoder, where the utility-oriented precoder and the cross-layer-modulated symbols can adjust the transmission power of each layer, and indirectly adjust the number of transmission layers, e.g., the layers with zero power are not sent. Therefore, the proposed CSI feedback not only directly adjusts the precoder, but also indirectly adjusts the number of transmission layers. Second, AI/ML-based control agents are also proposed that learn the mapping between estimated channel quality of signal to interference and noise ratio (SINR), and the transmission schemes that optimize the throughput of AI-native link, i.e.,
@@ III-B5 Cross-Module Optimization Process
Phase II then deactivates the above auxiliary loss with $\lambda=1$, focusing exclusively on BCE-driven end-to-end refinement to finally achieve the proposed cross-layer modulation and utility-oriented precoder. Notably, channel encoding and decoding modules are excluded during the training phase to ease the difficulty of training and are cascaded only in the inference phase to maintain compatibility with standard communication pipelines. It should be noted that the proposed bit-level bottlenecks bypass solution maintains the complete link differentiable for the above gradient-based optimization.
As for Phase III, the switching model $g_{\mathrm{la}}(\cdot)$ is trained using the loss of categorical cross-entropy, i.e.,
@@ III-B6 AI/ML Model Structure Implementation
The model structure supporting the proposed CSI feedback in III-B2 and III-B3 is detailed in Table III. The Transformer backbone for CSI feedback namely EVCsiNet-T han2024ai with embedding dimension of $N_{\mathrm{embedding}}=256$, $N_{\mathrm{head}}=4$ heads and $N_{\mathrm{transformer}}=6$ basic blocks is implemented for feature extraction, wherein the EVCsiNet-T is a common model structure used by 3GPP to evaluate the performance of CSI feedback. Specifically, the quantization and dequantization layers are replaced by Dense layers with normalization that directly output normalized modulation constellation symbols.
@@ III-C Theoretical analysis
In this subsection, we take the module of multi-layer modulation as an instance, shedding light on the potential benefits for performing cross-layer modulation with theoretical analysis. In our proposed transmission scheme, the bit-interleaved coded modulation (BICM) design principle is still inherited, where the performance of modulation can be measured by BICM capacity i2008bit
@@
where $b\in\{0,1\}$ is equiprobable bit, $m$ denotes modulation order, ${\bf{y}},{\bf{x}}\in{\text{C}}^{N_{\rm{layer}}}$ denote the received constellation and transmitted constellation, respectively. Notation $\mathcal{X}_{b}^{i}$ represents the set of symbols where the $i$-th bit equals $b$ and $\mathcal{X}$ denotes the set of all symbols. For traditional modulation method, both $\bf{y}$} and $\bf{x}$} are complex scalars. On the contrary, $\bf{y}$} and $\bf{x}$} can be complex vectors. The expectation operator shall jointly consider the high-dimensional distribution of transmitted symbols and noise, which is hard to be directly analyzed.
@@
Due to the intractability of the optimization on $l_{b}^{(i)}({\bf{y_{s}}})$ for all symbols, we turn to the analysis of upper bound of error probability for each symbol, which indirectly affects channel capacity. Reviewing the exponential component in (25), the minimization of probability is equivalent to maximizing the Euclidean distance between adjacent constellations, i.e. maximizing the minimum distance $d_{\rm{min}}$. Intuitively, a larger $d_{\rm{min}}$ for the given dimension can effectively reduce the error probability and contributes to improved channel capacity.
@@
where $r$ denotes the amplitude of the considered point, $V(\text{C}^{N_{\rm{layer}}})$ denotes the volume of the whole space. It can be expected that the shape of a sphere could maximize the space utilization efficiency, resulting in the averaged power of
@@
where the last approximation is obtained with $N!\simeq(N/e)^{N}$. These results are consistent with the ultimate shaping gain in [46]. To summarize, it can be seen that enlarging the modulation dimension can lead to an improved error probability and implicitly higher BICM capacity compared with legacy QAM constellation.
@@ IV Simulation Results
Downlink Carrier frequency 3.5GHz
@@
To demonstrate the effectiveness of the proposed cross-module optimization (CMO) scheme, we evaluate two training variants, i.e., CMO I and CMO II, against several 5G-compliant baselines under identical system constraints. CMO I exclusively undergoes phase I training until convergence, while CMO II sequentially executes phase I followed by phase II training. Both CMO I and CMO II share identical model implementations and total training iterations but employ distinct training processes. The conventional comparative scheme, denoted as 5G baselines, utilize standardized constellations for modulation and demodulation, including quadrature phase shift keying (QPSK), 16 quadrature amplitude modulation (QAM), 64QAM, and 256QAM, and supports a configurable number of transmission layers ranging in $N_{\mathrm{layer}}\in\{1,2,3,4\}$. The eTypeII codebook-based CSI feedback is adopted where eigenvectors that are fed back at the BS serve as the precoding matrix. To comprehensively assess baselines performance, three classical detection algorithms are employed. The LMMSE receiver aims to minimize the mean square error, thereby balancing interference suppression with noise enhancement. In contrast, the Zero-Forcing (ZF) receiver focuses solely on complete interference cancellation, which often leads to significant noise amplification. As a non-linear alternative, the K-Best receiver, a sub-optimal sphere decoding algorithm, provides a favorable performance-complexity trade-off by approximating the Maximum Likelihood (ML) performance. To focus solely on the downlink data transmission performance gains of the proposed CMO schemes, idealized uplink transmission conditions are assumed first. The CSI encoder and decoder components of both CMO I and CMO II retain the original EVCsiNet-T han2024ai architecture to ensure compatibility with CSI feedback overhead in bitstream form.
@@
As depicted in Fig. 4a to Fig. 4e, the BLER performance across varying downlink payload sizes is compared between the CMO II and CMO III approaches. The CMO III demonstrates consistent gains across all payload configurations, where the gain increases as the payload size grows. Notably, the CMO III scheme achieves more pronounced advantages under poorer uplink signal conditions. At an uplink SNR of -10 dB, taking BLER = 0.1 as a reference, the performance gain over the CMO II grows from approximately 1 dB in low-bit regions to nearly 5 dB in high-bit regions. In more severe scenarios with uplink SNR of -20 dB, the disparity widens significantly. That is, the gain escalates from around 4 dB at 2 bits per RE to up to 10 dB at 32 bits per RE under the same BLER threshold of 0.1. These results clearly demonstrate that the design enhancement for CSI feedback in CMO III effectively boosts communication performance in scenarios with limited uplink link quality. As illustrated in Fig. 4f, the throughput performance of both CMO II and CMO III under CCA is compared across varying signal conditions. Consistent with the trends observed in the BLER results, the CMO III exhibits more substantial throughput gains at lower uplink SNR, highlighting its robustness in UE power-constrained situation. At an uplink SNR of -10 dB, the CMO III achieves a throughput gain of approximately 22% over the CMO II when the downlink SNR is 10 dB. In contrast, under a more challenging uplink SNR of -20 dB, the gain increases to around 52% at the same downlink SNR level. These further explain the advantages of the proposed solutions.
@@
To rigorously validate the robustness of the proposed CMO III under more realistic and challenging conditions, we extend the evaluation to a scenario involving two BSs and two UEs coexisting. For downlink data transmission, the two BSs transmit signals over the same time-frequency resources, leading to mutual inter-BS interference. For uplink feedback, the two UEs also share identical resources for signal transmission, resulting in mutual inter-UE interference. In this experimental setup, the target-to-interference power ratio is set to 1/0.3 to mimic practical interference conditions. As illustrated in Fig. 6, under a payload size of 8 bits per RE, CMO III still maintains a significant performance gain over CMO II. Notably, the gain of CMO III over CMO II is more pronounced at a UL SNR of -20 dB than at -10 dB, aligning with the general trend observed in the interference-free scenario. Comparing Fig. 6 with the interference-free scenario in Fig. 4 reveals that CMO III exhibits minimal performance degradation, significantly smaller than that of CMO II. This result confirms that the end-to-end training under interference conditions enhances the overall interference tolerance at both the transmitter and receiver ends. More specifically, the enhanced CSI feedback mechanism in CMO III provides superior robustness against uplink interference, making it less affected by such interference than CMO II.